A 143143‑turn circular coil of radius 2.79 cm2.79 cm and negligible resistance is immersed in a uniform magnetic field that is perpendicular to the plane of the coil. The coil is connected to a 18.3 Ω18.3 Ω resistor to create a closed circuit. During a time interval of 0.159 s,0.159 s, the magnetic field strength decreases uniformly from 0.645 T0.645 T to zero. Find the energy, in millijoules, that is dissipated in the resistor during this time interval.

energy:

mJ
Answer #1

Part A:

Induced Emf in a coil is given by:

Emf = -N*d(phi)/dt

phi = Magnetic flux = B*A

Emf = -N*d(B*A)/dt = -N*A*dB/dt

Given that

N = 143 turns

A = pi*r^2

r = 2.79 cm = 0.0279 m

dB = 0 - 0.645 = -0.645 T

dt = 0.159 sec

So,

Emf = -143*pi*0.0279^2*(-0.645)/0.159

Emf = 1.42 V

Now Current through coil will be, Using Ohm's law:

I = Emf/R

R = resistance = 18.3 Ohm

I = 1.42/18.3

I = 0.0776 Amp

So, Power dissipated in the resistor will be

P = i^2*R = 0.0776^2*18.3

P = 0.1102 W = 110.2*10^-3 W

P = 110.2 mW

Now Energy required will be:

Energy = Power*time

Energy = (110.2 mW)*0.159 sec

Energy = 17.52 mJ

Please Upvote.